# From No-Regret to Strategically Robust Learning in Repeated Auctions

### v1

@@ Abstract
In this work, we show that this strategic robustness guarantee is not unique to agile OGD or to the first-price auction: any no-regret learning algorithm, when fed gradient feedback with respect to the quantile representation, is strategically robust, even if the auction format changes every round, provided the format satisfies allocation monotonicity and voluntary participation. In particular, the multiplicative weights update (MWU) algorithm simultaneously achieves the optimal regret guarantee and the best-known strategic robustness guarantee. At a technical level, our results are established via a simple relation that bridges Myerson’s auction theory and standard no-regret learning theory. This showcases the potential of translating standard regret guarantees into strategic robustness guarantees for specific games, without explicitly minimizing any form of swap regret.
@@ 1 Introduction
Consider a repeated single-item auction game over $T$ rounds between an auctioneer and $n$ bidders, where each bidder employs a no-regret learning algorithm [^1] Informally, an algorithm is no-regret if it performs at least as well as the best fixed strategy in hindsight., and the auctioneer, knowing bidders’ algorithms, strategically chooses the auction format in each round to maximize her cumulative revenue. This setting was first studied by Braverman et al. (2018), motivated by an empirical work of Nekipelov et al. (2015) showing that bidder behavior on Microsoft Bing is largely consistent with no-regret learning. This motivation has since been reinforced by the growing adoption of automated bidding agents that employ techniques from no-regret learning (Aggarwal et al., 2024).
Braverman et al. (2018) showed that many widely adopted no-regret learning algorithms, e.g., the multiplicative weights update (MWU) algorithm, are susceptible to strategic manipulation: a strategic auctioneer can choose auction formats dynamically to extract the full value of a bidder who adopts these algorithms, leaving the bidder with zero utility (a result extended to multiple bidders by Cai et al. (2023)), and moreover, she can obtain a higher average revenue per round than Myerson’s optimal auction by simply adjusting reserve prices over time in repeated first-price auctions. Their results have motivated the study of strategically robust no-regret learning algorithms that, when adopted by all bidders, cap the auctioneer’s average revenue per round at Myerson’s optimal revenue.
Most strategically robust no-regret algorithms designed so far explicitly minimize various forms of swap regret (Braverman et al., 2018, Mansour et al., 2022, Arunachaleswaran et al., 2025), which are stronger and more complex than standard regret. An exception is the recent work of Kumar et al. (2024), who showed that a natural no-regret learning algorithm––agile [^2] Briefly, agile OGD projects an iterate onto the feasible decision space at each update, whereas lazy OGD maintains an unconstrained iterate and projects only at decision time (Hazan, 2016, Section 5.4.1). online gradient descent (OGD)––is strategically robust in repeated first-price auctions with reserve prices. However, this result hinges on a subtle choice of the algorithm’s decision space: Kumar et al. (2024) apply agile OGD over what we call the quantile strategy space, which consists of quantile strategies (i.e., quantile representations of monotone bidding strategies; see Figure 1 for an illustration). Conceptually, their algorithm can be viewed as feeding the gradient of a bidder’s utility with respect to the quantile strategy as the reward vector to agile OGD.
@@
Figure 1: This figure illustrates a monotone (i.e., non-decreasing) bidding strategy that bids 0 for values in $0,v_{1}]$, bids $\frac{1}{2}$ for values in $(v_{1},v_{2}]$, and bids 11 otherwise. Suppose that $0,\frac{1}{2},1$ are the only possible bids, and let $F$ denote the CDF of the bidder’s value distribution. Kumar et al. (2024) represent a monotone bidding strategy using the quantiles of the value thresholds at which the bid increases ($F(v_{1})$ and $F(v_{2})$ in this case). We refer to this quantile-based representation as a quantile strategy. In this paper, we instead use the probability masses of the value intervals corresponding to distinct bids ($F(v_{1})$, $F(v_{2})-F(v_{1})$ and $1-F(v_{2})$ in this case) to represent the same quantile strategy, for a reason discussed in Section $1.$
This raises a natural question that inspired our work:
Does strategic robustness come from agile OGD itself, or the quantile strategy space it operates over?
Intuitively, agile OGD seems to be the source of strategic robustness: in the known examples of strategic manipulation (Braverman et al., 2018, Cai et al., 2023), the auctioneer exploits the lack of agility of algorithms such as MWU, whereas agile OGD naturally prevents such exploitation.
Contrary to this intuition, in this paper, we show that the quantile strategy space can be used to guarantee strategic robustness without relying on agile OGD. Specifically, we provide a meta-algorithm (Algorithm 1) that transforms any no-regret learning algorithm into a strategically robust one by feeding it the gradient of a bidder’s utility with respect to the quantile strategy as the reward vector. Moreover, the strategic robustness guarantee of our meta-algorithm holds even if the auctioneer changes the auction format every round, as long as the format satisfies allocation monotonicity and voluntary participation [^3] Informally, allocation monotonicity means that raising one’s bid cannot decrease the probability of winning the item, and our voluntary participation condition states that bidders can always ensure zero payment by bidding zero (see the formal statement in Assumption 2.4). These conditions are satisfied by most auctions deployed in practice..
Theorem 1.1 (Informal restatement of Theorem 4.3 and Proposition 4.2).
@@
In particular, MWU––the algorithm whose vulnerability to manipulation motivated the study of strategically robust no-regret learning––achieves the optimal regret and the best-known strategic robustness guarantee in our repeated auction setting, when combined with our meta-algorithm.
Corollary 1.2 (Informal restatement of Corollary 5.2).
If all bidders run Algorithm 1 instantiated with MWU, and the auction formats chosen by the auctioneer satisfy allocation monotonicity and voluntary participation, then the auctioneer’s expected cumulative revenue is at most $\textnormal{Mye}(\mathcal{D})\cdot T+O(n\cdot\sqrt{T\log(K)})$, where $\textnormal{Mye}(\mathcal{D})$ denotes Myerson’s optimal revenue, and $K$ is the number of possible bids. Moreover, every bidder’s regret is at most $O(\sqrt{T\log(K)})$.
For comparison, the strategic robustness guarantee of the algorithm in Braverman et al. (2018) has a $T^{3/4}$ dependence, while the guarantees of the algorithms in Mansour et al. (2022), Kumar et al. (2024), Arunachaleswaran et al. (2025) have various $K^{\Omega(1)}$ dependencies. However, we emphasize that these algorithms have their own merits beyond strategic robustness in our setting, and we refer interested readers to their respective papers for further details.
@@ 1.1 Technical overview
where $\bm{p}^{(t)}$ represents the bidder’s quantile strategy (see Figure 1 caption) at round $t$, and $u^{(t)}$ is the bidder’s utility function at round $t$. Using a potential function argument based on Eq. (1), they prove that agile OGD is strategically robust in the single-bidder case.
For the general setting with multiple bidders, their analysis follows the same plan as Braverman et al. (2018). Specifically, they first show that if a bidder adopts agile OGD, then in hindsight she can gain only negligible utility by swapping the bidding strategies associated with different value types across the entire time horizon (this property can be viewed as an online analog of incentive compatibility, or as a no-swap-regret property between value types). This, in turn, allows them to construct a global incentive-compatible auction that upper bounds the auctioneer’s average revenue per round. Their proof of the no-swap-regret property is based on another potential function argument inspired by their analysis of the single-bidder case, and it is carried out under a technical assumption that the PDFs of the bidders’ value distributions are bounded everywhere [^4] Our approach sidesteps this assumption..
@@
Then, we prove a simple fact––if each bidder $i\in[n]$ adopts our meta-algorithm, which feeds the gradient $\nabla u_{i}(\bm{p}_{i}^{(t)})$ as the reward vector to a no-regret learning algorithm $\mathcal{A}_{i}$, then the summation $\sum_{t\in[T]}\nabla u_{i}(\bm{p}_{i}^{(t)})^{\top}\bm{p}_{i}^{(t)}$ is upper bounded by algorithm $\mathcal{A}_{i}$’s regret. Thus, by summing both sides of Eq. (2) over $T$ rounds, we obtain that
@@
Finally, as mentioned in the caption of Figure 1, we use a different representation of quantile strategies in this paper. Instead of using the quantiles of the value thresholds at which the bid increases, we represent a quantile strategy by the probability masses of the value intervals that correspond to distinct bids. (Because of this choice, the relation in Eq. (2) takes a slightly different form in Lemma 4.7.) This transforms the decision space into a probability simplex, which enables MWU to recover its optimal regret guarantee and, together with our meta-algorithm, to achieve stronger strategic robustness than prior works.
@@ 1.2 Related work
There is a substantial literature on no-regret learning algorithms for repeated auctions (Feng et al., 2021, Balseiro et al., 2022, Zhang et al., 2022, Badanidiyuru et al., 2023, Aggarwal et al., 2025, Deng et al., 2025, Han et al., 2025). The study of the interaction between a strategic auctioneer and no-regret learning algorithms was initiated by Braverman et al. (2018), who showed that the auctioneer can extract the full value of a bidder adopting any mean-based no-regret algorithm, such as MWU. This result was later extended to multiple bidders by Cai et al. (2023). The same manipulation phenomenon has also been demonstrated for principal-agent problems (Guruganesh et al., 2024, Lin and Chen, 2024), first-price auctions between two bidders (Rubinstein and Zhao, 2024), and normal-form games (Deng et al., 2019b) (though computing a near-optimal manipulation strategy against MWU in normal-form games is $\mathsf{NP}$-hard (Assos et al., 2025)). These results have motivated research on designing no-regret learning algorithms that are strategically robust, for auctions (Braverman et al., 2018, Kumar et al., 2024) as well as for general Bayesian games and polytope games (Mansour et al., 2022, Arunachaleswaran et al., 2025).
@@
Most of these algorithms minimize various forms of swap regret. Specifically, Braverman et al. (2018) design an efficient algorithm––similar to the no-swap-regret algorithm in Blum and Mansour (2007)––to minimize a notion of swap regret between value types. Mansour et al. (2022) provide an (inefficient) algorithm, based on swap-to-standard-regret reductions (e.g., Stoltz and Lugosi (2005), Blum and Mansour (2007), Dagan et al. (2024), Peng and Rubinstein (2024)), to minimize a more sophisticated notion called polytope swap regret. Arunachaleswaran et al. (2025) design an efficient algorithm that minimizes a closely related notion called profile swap regret, through Blackwell’s approachability (Blackwell, 1956) and semi-separation (Daskalakis et al., 2025). In fact, minimizing appropriate notions of swap regret is both sufficient and necessary to achieve strategic robustness in general normal-form games (Deng et al., 2019b, Mansour et al., 2022) and in Bayesian and polytope games (Mansour et al., 2022, Rubinstein and Zhao, 2024, Arunachaleswaran et al., 2025).
@@
This makes the recent result of Kumar et al. (2024) particularly surprising: they show that agile OGD, operating over the quantile strategy space, is strategically robust in first-price auctions. Our work uncovers a general principle underlying this phenomenon—using the quantile strategy space, any no-regret learning algorithm can be transformed into a strategically robust one, even if the auctioneer is allowed to change the auction format over time, provided the format satisfies allocation monotonicity and voluntary participation.
@@
Furthermore, we note that there are also studies on strategic interactions between users who employ no-regret learning algorithms (e.g., Kolumbus and Nisan (2022a, b), Ananthakrishnan et al. (2024), Kolumbus et al. (2025)) and between heterogeneous agents (e.g., Easley et al. (2025)), as well as on other desirable properties of no-regret learning algorithms in strategic environments, such as Pareto optimality (Arunachaleswaran et al., 2024, 2025) and strategic robustness properties from the auctioneer’s perspective (e.g., Amin et al. (2013), Mohri and Munoz (2015), Liu et al. (2018), Deng et al. (2019a), Huh and Kirthevasan (2025)).
@@
Independently and concurrently, Yang Cai, Haipeng Luo, Chen-Yu Wei, and Weiqiang Zheng (private communication) have shown that any no-regret learning algorithm with gradient feedback under the quantile-based formulation of Kumar et al. (2024) is strategically robust in single-bidder repeated first-price auctions.
@@ 2.1 Single-item auctions
The auctioneer’s goal is to design an auction, which takes bidders’ bids as input and determines who gets the item and how much they pay, in order to maximize her expected revenue. Formally, an auction $\mathcal{M}:=(\mathbf{x},\mathbf{p})$ is specified by an allocation function $\mathbf{x}:[0,1]^{n}\to\Delta([n])$ (where $\Delta$ denotes the probability simplex) and a payment function $\mathbf{p}:[0,1]^{n}\to[0,1]^{n}$. Specifically, given a bid vector $\mathbf{b}=(b_{1},\dots,b_{n})$ as input, the allocation function $\mathbf{x}$ outputs a distribution $\mathbf{x}(\mathbf{b})$ over $[n]$, whose $i$-th coordinate $x_{i}(\mathbf{b})$ represents the probability that bidder $i$ receives the item, and the payment function $\mathbf{p}$ outputs a payment vector $\mathbf{p}(\mathbf{b})$, whose $i$-th coordinate $p_{i}(\mathbf{b})$ represents the amount which bidder $i$ pays the auctioneer.
An auction $\mathcal{M}=(\mathbf{x},\mathbf{p})$ is individual-rational (IR) if each bidder’s expected utility is guaranteed to be non-negative when bidding the true value, i.e., for all $i\in[n]$, for any $v_{i}\in[0,1]$ and $\mathbf{b}_{-i}:=(b_{1},\dots,b_{i-1},b_{i+1},\dots,b_{n})\in[0,1]^{n-1}$,
@@
where $(v_{i},\mathbf{b}_{-i}):=(b_{1},\dots,b_{i-1},v_{i},b_{i+1},\dots,b_{n})$. Moreover, the auction $\mathcal{M}$ is incentive-compatible (IC) if each bidder’s expected utility is maximized when bidding the true value, i.e., for all $i\in[n]$, for any $v_{i},b_{i}\in[0,1]$ and $\mathbf{b}_{-i}\in[0,1]^{n-1}$,

### v2

@@ Abstract
In this work, we show that this strategic robustness guarantee is not unique to agile OGD or to the first-price auction: any no-regret learning algorithm, when fed gradient feedback with respect to the quantile representation, is strategically robust, even if the auction format changes every round, provided the format satisfies allocation monotonicity and voluntary participation. In particular, the multiplicative weights update (MWU) algorithm simultaneously achieves the optimal regret guarantee and a strong strategic robustness guarantee in this auction setting. At a technical level, our results are established via a simple relation that bridges Myerson’s auction theory and standard no-regret learning theory.
@@ 1 Introduction
Consider a repeated single-item auction game over $T$ rounds between an auctioneer and $n$ bidders. In each round, each bidder employs a no-regret learning algorithm [^1] Informally, an algorithm is no-regret if it performs at least as well as the best fixed strategy in hindsight., while the auctioneer, knowing the bidders’ algorithms, strategically chooses the auction format to maximize her cumulative revenue. This setting was first studied by Braverman et al. (2018), motivated by an influential work of Nekipelov et al. (2015) showing that bidder behavior on Microsoft Bing is largely consistent with no-regret learning. This motivation has since been reinforced by the growing adoption of automated bidding agents that employ techniques from no-regret learning (Aggarwal et al., 2024).
Braverman et al. (2018) showed that many widely adopted no-regret learning algorithms, such as the multiplicative weights update (MWU) algorithm, are susceptible to strategic manipulation: a strategic auctioneer can dynamically choose auction formats to extract the full welfare of a bidder using these algorithms (a result later extended to multiple bidders by Cai et al. (2023)), leaving the bidder with zero utility; moreover, the auctioneer can obtain a higher average revenue per round than the revenue of Myerson’s optimal auction by simply adjusting reserve prices in repeated first-price auctions. These results motivated the study of strategically robust no-regret learning algorithms––algorithms that, when adopted by all bidders, cap the auctioneer’s average revenue per round at Myerson’s optimal revenue. Importantly, our focus is on the revenue cap in repeated single-item auctions (Braverman et al., 2018, Kumar et al., 2024); we do not address general non-manipulability in arbitrary normal-form or Bayesian games (Deng et al., 2019b, Mansour et al., 2022, Rubinstein and Zhao, 2024, Arunachaleswaran et al., 2025).
Most strategically robust (and general non-manipulable) no-regret algorithms designed so far explicitly minimize various notions of swap regret (Braverman et al., 2018, Mansour et al., 2022, Arunachaleswaran et al., 2025), which are stronger and more complex than standard regret. An exception is the recent work of Kumar et al. (2024), who showed that a natural no-regret learning algorithm––agile [^2] Briefly, agile OGD projects an iterate onto the feasible decision space at each update, whereas lazy OGD maintains an unconstrained iterate and projects only at decision time (Hazan, 2016, Section 5.4.1). online gradient descent (OGD)––is strategically robust in repeated first-price auctions with reserve prices. A key aspect of their result is the choice of decision space: Kumar et al. (2024) apply agile OGD over what we call the quantile strategy space, which consists of quantile representations of monotone bidding strategies (see Figure 1). Conceptually, their algorithm can be viewed as feeding the gradient of a bidder’s utility with respect to the quantile strategy as the reward vector to agile OGD.
@@
Figure 1: This figure illustrates a monotone (i.e., non-decreasing) bidding strategy that bids 0 for values in $0,v_{1}]$, bids $\frac{1}{2}$ for values in $(v_{1},v_{2}]$, and bids 11 otherwise. Suppose that $0,\frac{1}{2},1$ are the only possible bids, and let $F$ denote the CDF of the bidder’s value distribution. Kumar et al. (2024) represent a monotone bidding strategy using the upper-tail probabilities of the value thresholds at which the bid increases ($1-F(v_{1})$ and $1-F(v_{2})$ in this case). We refer to this quantile-based representation as a quantile strategy. In this paper, we instead use the probability masses of the value intervals corresponding to distinct bids ($F(v_{1})$, $F(v_{2})-F(v_{1})$ and $1-F(v_{2})$ in this case) to represent the same quantile strategy, for a reason discussed in Section $1.$
This observation raises a natural structural question that motivates our work:
What role does the quantile strategy space itself play in enabling strategic robustness, independent of the specific learning dynamics used?
In this paper, we show that the quantile strategy space itself plays an important role: strategic robustness can be guaranteed without relying on agile OGD. Specifically, we provide a meta-algorithm (Algorithm 1) that transforms any no-regret learning algorithm into a strategically robust one by feeding it the gradient of a bidder’s utility with respect to the quantile strategy as the reward vector. The resulting strategic robustness guarantee holds even if the auctioneer changes the auction format in every round, provided the format satisfies allocation monotonicity and voluntary participation [^3] Informally, allocation monotonicity means that raising one’s bid cannot decrease the probability of winning the item, and our voluntary participation condition states that bidders can always ensure zero payment by bidding zero (see Assumption 2.4). These conditions are satisfied by most auction formats deployed in practice..
Theorem 1.1 (Informal statement of Theorem 4.3 and Proposition 4.2).
@@
In particular, when combined with our meta-algorithm, MWU––despite its known vulnerability to manipulation––achieves optimal regret and strong strategic robustness in our auction setting. The strategic robustness guarantee does not stem from any new property of MWU itself, but from the structural transformation induced by our meta-algorithm.
Corollary 1.2 (Informal statement of Corollary 5.2).
If all bidders run Algorithm 1 instantiated with MWU, and the auction formats chosen by the auctioneer satisfy allocation monotonicity and voluntary participation, then the auctioneer’s expected cumulative revenue is at most $\textnormal{Mye}(\mathcal{D})\cdot T+O(n\cdot\sqrt{T\log(K)})$, where $K$ is the number of possible bids. Moreover, each bidder’s regret is at most $O(\sqrt{T\log(K)})$.
For comparison, the strategic robustness guarantee of the algorithm in Braverman et al. (2018) has a $T^{3/4}$ dependence, while the guarantees of the algorithms in Mansour et al. (2022), Kumar et al. (2024), Arunachaleswaran et al. (2025), when applied to the auction setting, exhibit various $K^{\Omega(1)}$ dependencies. However, these algorithms address broader or stronger objectives––such as general non-manipulability in Bayesian and polytope games or incentive compatibility––beyond the revenue cap studied here.
@@ 1.1 Technical overview
where $\bm{p}^{(t)}$ represents the bidder’s quantile strategy (see Figure 1 caption) at round $t$, and $u^{(t)}$ is the bidder’s utility function at round $t$. Using a potential function argument based on Eq. (1), they prove that agile OGD, when used to update the quantile strategy, is strategically robust in the single-bidder case.
For the general setting with multiple bidders, their analysis follows the same plan as Braverman et al. (2018). Specifically, they first show that if a bidder adopts agile OGD (in a manner similar to the single-bidder case), then in hindsight she can gain only negligible utility by swapping the bidding strategies associated with different value types across the entire time horizon (this property can be viewed as an online analog of incentive compatibility, or as a no-swap-regret property between value types). This, in turn, allows them to construct a global incentive-compatible auction that upper bounds the auctioneer’s average revenue per round. Their proof of the no-swap-regret property is based on another potential function argument inspired by their analysis of the single-bidder case, and it is carried out under a technical assumption that the PDFs of the bidders’ value distributions are bounded everywhere [^4] Our approach sidesteps this assumption..
@@
Then, we prove a simple but crucial fact––if each bidder $i\in[n]$ adopts our meta-algorithm, which feeds the gradient $\nabla u_{i}^{(t)}(\bm{p}_{i}^{(t)})$ as the reward vector to a no-regret learning algorithm $\mathcal{A}_{i}$, then the summation $\sum_{t\in[T]}\nabla u_{i}^{(t)}(\bm{p}_{i}^{(t)})^{\top}\bm{p}_{i}^{(t)}$ is upper bounded by algorithm $\mathcal{A}_{i}$’s regret. Thus, by summing both sides of Eq. (2) over $T$ rounds, we obtain that
@@
Finally, as mentioned in the caption of Figure 1, we use a different representation of quantile strategies in this paper. Instead of using the upper-tail probabilities of the value thresholds at which the bid increases, we represent a quantile strategy by the probability masses of the value intervals that correspond to distinct bids. (Because of this choice, the relation in Eq. (2) takes a slightly different form in Lemma 4.7.) This transforms the decision space into a probability simplex, which enables MWU to recover its optimal regret guarantee and, together with our meta-algorithm, to achieve strong strategic robustness.
@@ 1.2 Related work
There is a substantial literature on no-regret learning algorithms for repeated auctions (Feng et al., 2021, Balseiro et al., 2022, Zhang et al., 2022, Badanidiyuru et al., 2023, Aggarwal et al., 2025, Deng et al., 2025, Han et al., 2025). The study of the interaction between a strategic auctioneer and no-regret learning algorithms was initiated by Braverman et al. (2018), who showed that the auctioneer can extract the full welfare of a bidder adopting any mean-based no-regret algorithm, such as MWU. This result was later extended to multiple bidders by Cai et al. (2023). Similar manipulation phenomena have also been demonstrated in related settings, including principal-agent problems (Guruganesh et al., 2024, Lin and Chen, 2024), first-price auctions between two bidders (Rubinstein and Zhao, 2024), and normal-form games (Deng et al., 2019b). In the latter case, computing a near-optimal manipulation strategy against MWU is known to be $\mathsf{NP}$-hard (Assos et al., 2025). Together, these results motivate the design of no-regret learning algorithms that are non-manipulable in general normal-form, Bayesian, and polytope games (Deng et al., 2019b, Mansour et al., 2022, Arunachaleswaran et al., 2025), or strategically robust in auction environments (Braverman et al., 2018, Kumar et al., 2024).
@@
A prominent line of work achieves general non-manipulability (or strategic robustness in auction environments) by minimizing various notions of swap regret. Braverman et al. (2018) design an efficient algorithm for the auction setting with bandit feedback—inspired by the no-swap-regret algorithm of Blum and Mansour (2007)—to minimize a notion of swap regret between value types. For more general Bayesian and polytope games, Mansour et al. (2022) provide an (exponential-time) algorithm based on the no-swap-regret algorithm of Blum and Mansour (2007) to minimize a notion called polytope swap regret, and Arunachaleswaran et al. (2025) design an efficient algorithm to minimize profile swap regret, using Blackwell’s approachability (Blackwell, 1956) and semi-separation (Daskalakis et al., 2025). In fact, minimizing appropriate notions of swap regret is both sufficient and necessary to achieve non-manipulability in general normal-form games (Deng et al., 2019b, Mansour et al., 2022) as well as in Bayesian and polytope games (Mansour et al., 2022, Rubinstein and Zhao, 2024, Arunachaleswaran et al., 2025).
@@
Our work is complementary to this literature. Rather than studying general non-manipulability for general games, we focus on a structured auction environment where auction-specific properties can be leveraged to obtain revenue caps. In this context, the recent result of Kumar et al. (2024) highlights a distinct phenomenon: they show that agile OGD, when operating over the quantile strategy space, is strategically robust in repeated first-price auctions with reserve prices. Our results identify a general principle underlying this phenomenon: by learning in the quantile strategy space, any no-regret learning algorithm can be transformed into a strategically robust one for repeated single-item auctions, even when the auctioneer is allowed to change the auction format over time, provided the format satisfies allocation monotonicity and voluntary participation.
@@
Finally, we note that there are studies on strategic interactions between users who employ no-regret learning algorithms (e.g., Kolumbus and Nisan (2022a, b), Ananthakrishnan et al. (2024), Kolumbus et al. (2025)) and between heterogeneous agents (Easley et al., 2025). Other works study additional desirable properties of no-regret learning in strategic environments, such as Pareto optimality (Arunachaleswaran et al., 2024, 2025) and strategic robustness guarantees from the auctioneer’s perspective (e.g., Amin et al. (2013), Mohri and Munoz (2015), Liu et al. (2018), Deng et al. (2019a), Huh and Kirthevasan (2025)).
@@
Independently and concurrently, Yang Cai, Haipeng Luo, Chen-Yu Wei, and Weiqiang Zheng (private communication) have shown that any no-regret learning algorithm with gradient feedback under the quantile-based formulation is strategically robust in single-bidder repeated first-price auctions.
@@ 2.1 Single-item auctions
The auctioneer’s goal is to design an auction, which takes bidders’ bids as input and determines who gets the item and how much they pay, in order to maximize her expected revenue. Formally, an auction $\mathcal{M}:=(\mathbf{x},\mathbf{p})$ is specified by an allocation function $\mathbf{x}:[0,1]^{n}\to\mathcal{X}([n])$ (where $\mathcal{X}([n]):=\left\{\mathbf{y}\in\mathbb{R}_{\geq 0}^{n}\;\middle|\;\sum_{i=1}^{n}y_{i}\leq 1\right\}$) and a payment function $\mathbf{p}:[0,1]^{n}\to[0,1]^{n}$. Specifically, given a bid vector $\mathbf{b}=(b_{1},\dots,b_{n})$ as input, the $i$-th coordinate $x_{i}(\mathbf{b})$ of the output of the allocation function $\mathbf{x}$ is the probability that bidder $i$ receives the item, and the $i$-th coordinate $p_{i}(\mathbf{b})$ of the output of the payment function $\mathbf{p}$ is the amount which bidder $i$ pays the auctioneer.
An auction $\mathcal{M}=(\mathbf{x},\mathbf{p})$ is individual-rational (IR) if each bidder’s expected utility is guaranteed to be non-negative when bidding her true value, i.e., for all $i\in[n]$, for any $v_{i}\in[0,1]$ and $\mathbf{b}_{-i}:=(b_{1},\dots,b_{i-1},b_{i+1},\dots,b_{n})\in[0,1]^{n-1}$,
@@
where $(v_{i},\mathbf{b}_{-i}):=(b_{1},\dots,b_{i-1},v_{i},b_{i+1},\dots,b_{n})$. Moreover, the auction $\mathcal{M}$ is incentive-compatible (IC) if each bidder’s expected utility is maximized when bidding her true value, i.e., for all $i\in[n]$, for any $v_{i},b_{i}\in[0,1]$ and $\mathbf{b}_{-i}\in[0,1]^{n-1}$,